Find the equations of tangent line and normal lines of parabola y= 2x+ x squared at point (1,3).

Find all the values of x at which the tangent line to the curve y=1/ x + 4 passes through the origin.
1
Expert's answer
2019-08-12T09:30:25-0400

Find the equations of tangent line and normal lines of parabola y= 2x+ x squared at point (1,3).


"y=2x+x^2, \space \space \space (1,3)""y'=2+2x, \space \space \space y'(1)=2+2*1=4."

Tangent line:

"y=y'_0(x-x_0)+y_0"

"y=4(x-1)+3"

"y=4x-1."

Normal line:

"y=-\frac{1}{y'_0}(x-x_0)+y_0"

"y=-\frac{1}{4}(x-1)+3"

"y=\frac{-x+13}{4}."


Find all the values of x at which the tangent line to the curve y=1/ x + 4 passes through the origin.


"y=\frac 1 {x+4}, \space\space y'=-\frac 1 {(x+4)^2}"

"y=y'_0(x-x_0)+y_0"

"y=y'_0x+(-y'_0x_0+y_0)"

line passes through the origin => "-y'_0x_0+y_0=0"


"-\frac 1 {(x_0+4)^2}*(-x_0)+\frac 1 {x_0+4}=0"

"\frac {x_0} {(x_0+4)^2}+\frac {x_0+4} {(x_0+4)^2}=0"

"\frac {2x_0+4} {(x_0+4)^2}=0"

"2x_0+4=0 \space and \space x_0+4\not=0"

"x_0=-2."

Answer:

1) Tangent line "y=4x-1," normal line "y=\frac{-x+13}{4}."

2) x=-2.